OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year?

OpenStudy (mathmate):

Use the compound interest formula: FV=PV*(1+r)^t or FV/PV=2=(1+r)^t Here FV/PV=2 (double) r=6% p.a., so 3% for 6 months. Take log on each side and divide to solve for t: log(2)=tlog(1+r/2) t=log(2)/log(1.03)=23.45 periods of 6 months, or 11.72 years approx. You can check your answer using the rule of 72, which is 72/interest rate=number of years to double. 72/6=12. Since it is compounded every 6 months, it takes slightly less time. If it were compounded yearly, it would have taken log(2)/log(1.06)=11.9 years.

OpenStudy (anonymous):

thank you sooo much (:

OpenStudy (mathmate):

You're welcome!